BestBoiKirby:

i need halp with more khan academy. (its a different topic than what i just posted like 15 mins ago, it just looks similar)

BestBoiKirby:

1 attachment
Acolgrove:

the dreaded khan academy

BestBoiKirby:

who should i ping for this?

BestBoiKirby:

@umm you had human calculator so i ping you

Acolgrove:

um yeah i am so bad at math

BestBoiKirby:

@darkknight

BestBoiKirby:

im pinging people with math titles

Acolgrove:

i would ping mira shes good at math i think

BestBoiKirby:

@MiraAngel

BestBoiKirby:

@silvernight269

BestBoiKirby:

@justus @Pixel

BestBoiKirby:

there i pinged all the peeps with math titles i need halp D:

Acolgrove:

yay silver lol

BestBoiKirby:

1 attachment
BestBoiKirby:

SILVERRRR do you know how to do this?

silvernight269:

kirby son theres a thing called watch video. Its khan academy

BestBoiKirby:

but i dont get it D:

silvernight269:

gimme a few ill try it

silvernight269:

what is the point slope formula

BestBoiKirby:

absolutely no idea

Acolgrove:

lol silver

MiraAngel:

y-y₁=m(x-x₁)

silvernight269:

y-y1=m(x-x1)

silvernight269:

basically what mira said.

BestBoiKirby:

it says that is has a wrong variable letter

MiraAngel:

m is slope. Do you know how to find slope?

BestBoiKirby:

i know nothing abt this D:

MiraAngel:

Okay well just remember rise over run... y2-y1/x2-x1

MiraAngel:

Geez! I'm sorry Silver! This is your question! You answered first. That was rude. Go ahead.

BestBoiKirby:

ok, well i also put in the thing and it said there an unknown variable

BestBoiKirby:

hang on lemme check something real quick

MiraAngel:

Okay

BestBoiKirby:

yeah i do not understand any of this

darkknight:

Okay, you know how to take the slope?

BestBoiKirby:

it said the answer is y-(-1)=3/4(x-1)

BestBoiKirby:

and no

BestBoiKirby:

hang on i gotta restart it now that it gave me the hints, it counts it as wrong

BestBoiKirby:

1 attachment
BestBoiKirby:

heres the new problem

darkknight:

Alright, same steps are needed for this problem. Basically you need to find the slope which I am going to call m \[m = (y2 - y1)/(x2-x1)\] (the 1's and the 2's above supposed to be in subscript

MiraAngel:

Okay. You need to find slope. y2-y1/x2-x1 (x1, y1) (x2, y2) (1, -1) (5, 2)

BestBoiKirby:

Sooooo

BestBoiKirby:

6-1=5

BestBoiKirby:

-2-1=-3?

darkknight:

no, you have to take the slope, so first you take the change in the y values

BestBoiKirby:

i did that

darkknight:

What is 2 minus -1?

BestBoiKirby:

ohhh

BestBoiKirby:

wait -2 and 1 are the x values tho

BestBoiKirby:

6 and 1 are the y values

MiraAngel:

(-2, 6) (1, 1) m= y2-y1/x2-x1 1-6/1-(-2) -5/3

darkknight:

Nope other way around, y values are after x values written like this (x,y)

BestBoiKirby:

exactly

BestBoiKirby:

and look at the picture

BestBoiKirby:

1 attachment
BestBoiKirby:

i thought i had it for a second and i didnt and i got it wrong.... OK IMMA REPOST ONE LAST TIME

BestBoiKirby:

1 attachment
BestBoiKirby:

there it is

MiraAngel:

y-y₁=m(x-x₁) y-6= (-5/3) x+2

BestBoiKirby:

mira thank u but i already got it wrong, i thought i had it but i didnt, heres the new one

1 attachment
MiraAngel:

kk First slope. Can you remind me what the formula for that is?

darkknight:

bro @BestBoiKirby don't try doing it without knowing what you are doing. Do you know how to calculate slope?

BestBoiKirby:

so its x-x over y-y right?

BestBoiKirby:

wait no y-y over x-x

BestBoiKirby:

with x on bottom and y on top

BestBoiKirby:

right?

MiraAngel:

First, write out what you know... (-1, 6) and (1, 5) y-6=..... y-y1=m(x-x1) Slope: m= y2-y1/x2-x1 5-6/1-(-1) m= -1/2 y-6= (-1/2) x-(-1) y-6=(-1/2) x+1

BestBoiKirby:

huh?

MiraAngel:

I did the whole problem... There's all my work

BestBoiKirby:

OHHH i thought that was just one step

BestBoiKirby:

hang on so thats how u do it for every dingle one?

madisonlacroix:

...

MiraAngel:

The main thing you need to focus on and find is slope, which is M. It gives you everything else and you can input it into the equation easily.

MiraAngel:

Yeah

BestBoiKirby:

okie, ill post on here again if i have anymore issues, thank you!

MiraAngel:

Okay! Great! Good luck! :)

BestBoiKirby:

it says that was wrong

BestBoiKirby:

i put this one there y-6=(-1/2) x+1

BestBoiKirby:

y-6= (-1/2) x-(-1) but it said the step before was the correct answer

MiraAngel:

Oh. Okay. Well those two are technically the same thing. I just reduced it further. I would recommend emailing or talking to your teacher. Just let them know that you changed the minus negative into a plus type of problem. Two negatives equal a positive.... I just reduced it. The problem didn't

BestBoiKirby:

oh okie

darkknight:

actually u did something wrong here y-6=(-1/2) x+1 just means that -1/2 is being multiplied to x, not x+1, YOu need to write it like this y-6=(-1/2) (x+1)

MiraAngel:

^^^^ That makes sense too. lol That's why having a second/ third mind available is always safe. ^_^